7x+8y=25
9x+10y=35 using substitution or elimination

Found 2 solutions by josgarithmetic, MathTherapy:
Answer by josgarithmetic(39618)

Substitution: Pick an equation; either equation. Solve for one of the variables in terms of the other. Substitute this expression in the other equation and solve for the value of the single variable.
...

Answer by MathTherapy(10552)


7x+8y=25
9x+10y=35 using substitution or elimination

7x + 8y = 25 ------- eq (i)
9x + 10y = 35 ------- eq (ii)
- 35x - 40y = - 125 ------- Multiplying eq (i) by - 5 ------ eq (iii)
36x + 40y = 140 ------- Multiplying eq (ii) by 4 ------- eq (iv)
x = 15 ------ Adding eqs (iii) & (iv)
Now substitute 15 for x in eq (i) or eq (ii), or any of the other equations to determine the value of y